OpenStudy (anonymous):

How do you c hange 0.05 to a fraction/mixed number?

OpenStudy (anonymous):

do i just knock off the 0 and put 5/100?

OpenStudy (maheshmeghwal9):

\[\huge{\color{red}{0.05=\frac{5}{100}=\frac{1}{20}.}}\]

OpenStudy (maheshmeghwal9):

ya @etina02

OpenStudy (anonymous):

count how many places behind the decimal place the last digit is here, it's two then you put the number, without the decimal, over 10^n, where n is the number of places behind the decimal so you get 5/10^2 = 5/100 then simplify

OpenStudy (anonymous):

ok...got it....thank you.

OpenStudy (maheshmeghwal9):

yw:)

OpenStudy (anonymous):

are you arabic?

OpenStudy (anonymous):

indian i'm guessing

OpenStudy (maheshmeghwal9):

ya good catch @cerezas

OpenStudy (anonymous):

palestinian....from jordan or egypt?

OpenStudy (maheshmeghwal9):

but for what reason u said me arabic? @etina02

OpenStudy (anonymous):

your name

OpenStudy (anonymous):

indian, from india @maheshmeghwal9 probs your profile pic of john abraham XD

OpenStudy (maheshmeghwal9):

ya:) so u too indian @cerezas

OpenStudy (maheshmeghwal9):

my name-> Mahesh.

OpenStudy (anonymous):

yup

OpenStudy (maheshmeghwal9):

hmm....

OpenStudy (anonymous):

i am native american from the west coast of united states.

OpenStudy (maheshmeghwal9):

k! so nice to meet an american:)

OpenStudy (anonymous):

that's actually really awesome @etina02 where on the west coast? i'm headed to cali for college

OpenStudy (anonymous):

oh really????? you have meant one before..............but thanks............so you love math?

OpenStudy (anonymous):

...which of us?

OpenStudy (maheshmeghwal9):

ya?

OpenStudy (anonymous):

mahesh......but you too....cerezas.....

OpenStudy (maheshmeghwal9):

ya of course I love maths but it does not mean that I m a 100% scorer in maths.

OpenStudy (anonymous):

my boyfriend's part cherokee, so yeah and i really enjoy abstract math - i better, my college require five terms of math

OpenStudy (maheshmeghwal9):

sometimes i also get confused:/

OpenStudy (anonymous):

my name's ellora, btw it's weird being cerezas all the time

OpenStudy (anonymous):

your picture shows you not too confused but confidence MAHESH.......ha ha ha....my name is Tina...your name ellora is pretty.

OpenStudy (anonymous):

thank you :) i like it @etina02 it's not him, its a movie star in india :P

OpenStudy (maheshmeghwal9):

ya

OpenStudy (anonymous):

by the way......my email is etina02@yahoo.com..........both of you email me.......Mahesh...and Ellora.

OpenStudy (anonymous):

library is closing.............got to go.......will i hope i can contact you both with math problems and chatting on email.....bye......

OpenStudy (anonymous):

sure :) mine is sarkar.ellora@gmail.com

OpenStudy (maheshmeghwal9):

maheshmeghwal9@gmail.com

OpenStudy (anonymous):

bye

OpenStudy (maheshmeghwal9):

bye

OpenStudy (anonymous):

bye

OpenStudy (maheshmeghwal9):

@etina do not go on the wall pictures of anyone Sometimes they aren't real:)

OpenStudy (maheshmeghwal9):

as my

OpenStudy (maheshmeghwal9):

hope @cerezas have real:D

OpenStudy (anonymous):

yeah that's me

OpenStudy (maheshmeghwal9):

looking nice xD

OpenStudy (anonymous):

haha thanks :)

OpenStudy (maheshmeghwal9):

yw:) k! bye:)

OpenStudy (anonymous):

bye :)